The exhaust gas leaving an oil engine is passed through a heat exchanger which
consist of tubes surrounding by water. The gas enters at 327 oC and leaves at 193
oC. The water enters the exchanger at 17 and leaves at 56 oC. If the mass flow rate
of water is 2.54kg/min. Determine the mass flow rate of gas, take cp of the gas is
1.04 kJ/kg.
1
Expert's answer
2020-03-05T07:58:30-0500

"Q=cm(t_2-t_1)"


So, we have


"c_{pw}\dot{m}_w(t_{2w}-t_{1w})=c_{pg}\dot{m}_g(t_{1g}-t_{2g})"


"\dot{m}_g=\frac{c_{pw}\dot{m}_w(t_{2w}-t_{1w})}{c_{pg}(t_{1g}-t_{2g})}=
\frac{4187\cdot 2.54\cdot (56-17)}{1040\cdot(327-193)}=2.98kg/min" Answer